having a difficult time with finding the answer to a - 8 = 23? how do i solve a? and 13 + z = 6 how do I solve for z

Found 2 solutions by kingster64, tutorcecilia:
Answer by kingster64(2)

For the first one you will need to isolate a by adding 8 to both sides. Therefore, a will equal 23+8 = 31
a=31
a-8 = 31
a+ 8 = 8 ( the negative 8 plus positive will cancel each other out and leave a)
a=31

Answer by tutorcecilia(2152)

a - 8 = 23
a-8+8=23+8
a=31
Checking:
Let a=31
(31)-8=23
23=23
.
.
13 + z = 6
13-13+z=6-13
z=-7
Checking:
Let z=-7
13+(-7)=6
6=6